OpenStudy (anonymous):

What is the probability of rolling a total of 7 in two rolls of a standard die if you get an even number on the first roll?

OpenStudy (anonymous):

first roll even, so 2, 4 or 6 to make up seven, second roll can only be 5, 3 or 1 p(first even) = 3/6=0.5 p(second odd) = 3/6= 0.5 so p(even first, odd second) = 0.5 times 0.5 = 0.25, or 25%

OpenStudy (anonymous):

wrongs its 1 /6

OpenStudy (anonymous):

you have to use conditional probability i figured out